OpenStudy (anonymous):

how to prove tan^(-1) x +tan(^-1) y= tan^(-1) (x+y)/(1-xy)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \tan^{-1}(x)+\tan^{-1}(y)=\frac{\tan^{-1}(x+y) }{1-xy}}\) like this?

OpenStudy (anonymous):

the answer is \[\tan ^{-1} \frac{ x+y }{ 1-xy }\]

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \tan^{-1}(x)+\tan^{-1}(y)=\tan^{-1}\left(\frac{x+y }{1-xy}\right)}\) Like that you mean?

OpenStudy (anonymous):

yup

myininaya (myininaya):

I think you mistranslated @mathmale .

myininaya (myininaya):

try taking tan( ) of both sides @eunna

OpenStudy (anonymous):

then?

myininaya (myininaya):

Do you know sum rule for tan?

OpenStudy (anonymous):

no

myininaya (myininaya):

What identities are you allowed to use then?

OpenStudy (anonymous):

ohh the one that tan(a+b)

myininaya (myininaya):

yes

myininaya (myininaya):

a+b is read as the sum of a and b

OpenStudy (anonymous):

after apply the sum rule, how can i make it inverse?

myininaya (myininaya):

do you know what tan(arctan(u)) equals?

OpenStudy (anonymous):

equal to u

myininaya (myininaya):

right

OpenStudy (anonymous):

so i just need to multiply with tan^-1

OpenStudy (anonymous):

wont it eliminate the tan

myininaya (myininaya):

multiply?

myininaya (myininaya):

Have you taken tan( ) of both sides and applied sum rule yet for tan? and then use that tan(arctan(u))=u?

myininaya (myininaya):

If you are having trouble understanding this way, you could try doing the following: Let arctan(x)=u and arctan(y)=v and so tan(u)=x and tan(v)=y... Plug tan(u)=x and tan(v)=y into the right hand side... You will still have to use sum rule for tan...

myininaya (myininaya):

for example you know \[\frac{\tan(u)+\tan(v)}{1-\tan(u)\tan(v)}=?\]

OpenStudy (anonymous):

i got it, thank you so much

myininaya (myininaya):

np